David Z. Maze

92 Packard Ave. #2 Somerville, MA 02144 (617) 625-4297 dmaze@mit.edu

Objective

To find a permanent position in the field of Electrical Engineering and Computer Science.

Education

Massachusetts Institute of Technology.......................................................................................................................Cambridge, MA Recieved BS in Electrical Engineering and Computer Science in June 2000; will recieve M.Eng. in September 2001 Classes taken include Structure and Interpretation of Computer Programs, Differential Equations, LEGO Robot Design Competition, Computation Structures, Signals and Systems, Computer System Engineering, Laboratory in Software Engineering, Artificial Intelligence, Computer Language Engineering, Computer System Architecture, Multithreaded Parallelism: Languages and Compilers, Computer Graphics, Introductory Digital Design Laboratory, Network Optimization.

Experience

Akamai Technologies .....................................................................................................................................................Cambridge, MA Software Intern October 1999 — April 2001 Enhanced build infrastructure for source code. Developed Web-based administration tools for Perforce source control management system. Assisted in deployment of automated testing infrastructure.

Altera Corporation................................................................................................................................................................San Jose, CA Software Intern May 1999 — August 1999 Added an optimization stage to a hardware description language compiler for programmable logic devices.

MIT Information Systems.............................................................................................................................................Cambridge, MA Programmer June 1997 — May 1999 Developed World Wide Web-based front end to existing housing lottery system. Modified xscreensaver for use in the Athena Computing Environment. Also worked on infrastructure for a job-queueing system.

MIT Media Lab................................................................................................................................................................Cambridge, MA Programmer January 1997 — May 1997 Worked on Canard, a project to explore a community’s interactions using two-way alphanumeric pagers. Developed a system to send notification of incoming electronic mail to pagers.

Proficiencies

Over six years of experience with UNIX, and familiarity with MacOS, Windows, Linux, Solaris, IRIX. Programming experience includes C, C++, Java, Perl, PostScript, Scheme, Haskell.

Interests and Background

MIT Lecture Series Committee ..............................................................................................................September 1996 — May 2000 The Lecture Series Committee presents guest speakers and feature films to the MIT community. Work has included ticket sales and collection, supervision of other workers, screening of films, and movie selection. Positions held include Friday Worker, Night Worker, Night Subdirector, and Night Director. Member of LSC Executive Committee from April 1998 to December 1999.

Student Information Processing Board .....................................................................................................September 1999 — present SIPB provides software, support, and computer services to users of MIT’s Athena Computing Environment. Member-at-large of the SIPB Executive Committee from February 2000 to September 2000. Secretary from September 2000 to December 2000.

Other Activities Other extracurricular activities include the Living Group Council, the MIT Assassins’ Guild, and Alpha Phi Omega, a coed service fraternity. High school activities included speech and debate team and section editor for school newspaper.